ou have diluted 2.00mL of 7.30mM iodine solution to a total volume of 8.00mL with acetone, acid and water. It has taken 5.01 minutes for all of the color to disappear. What is the rate of this reaction in units of mM/min
Answer #1

Volume of Iodine solution V1 = 2.00ml

Concentration of Iodine solution M1 = 7.30mM

Volume after dilution V2 = 8.00ml

Therefore concentration of the dilute solution M2 = M1V1/V2 = 7.30*2/8 = 1.825mM

Time taken for the color to disappear = 5.01 minutes

Therefore the rate of reaction = 1.825mM/5.01 minutes = 0.3643mM/min